# Entropy-driven entanglement forging

### v1

@@ Abstract
Simulating a physical system with variational quantum algorithms is a well-studied approach but challenging to implement in current devices due to demands in qubit number and circuit depth. We show how limited knowledge of the system, namely the entropy of its subsystems or its entanglement structure, can be used to reduce the cost of these algorithms with entanglement forging. To do so, we simulate a Fermi-Hubbard one-dimensional chain with a parametrized hopping term, as well as atomic nuclei 28Ne and 60Ti with the nuclear shell model. Using an adaptive variational quantum eigensolver we find significant reductions in both the maximum number of qubits (up to one fourth) and the amount of two-qubit gates (over an order of magnitude) required in the quantum circuits. Our findings indicate that our method, entropy-driven entanglement forging, can be used to adjust quantum simulations to the limitations of current noisy intermediate-scale quantum devices.
@@ I Introduction
In modern quantum devices, the number of available qubits and low-error quantum gates imposes a strong limitation in the accuracy of the final results not only in fault-tolerant schemes but also in current noisy simulations. Since quantum algorithms are applied to complex quantum many-body problems [1, 2], including quantum chemistry [3, 4, 5], condensed matter  [6, 7] and nuclear physics [8, 9, 10], this constraint has been the source of many techniques that aim to simulate large systems with smaller but equivalent ones needing fewer quantum resources. Some encoding methods aim to reduce the dimension of the input data, like quantum autoencoding [11] or other physically-inspired frameworks [12]. On the other hand, other approaches focus directly on circuits to run equivalent simulations using fewer qubits. The best-known of them, circuit knitting [13, 14], takes advantage of circuits with sparsely-connected subsections in order to break them apart into smaller ones. More recently, an alternative approach has also been explored: one can train small circuits to prepare local states, and then recover the global solution with post-processing. Both types of techniques are commonly used on variational quantum algorithms (VQA) [1, 2].
Entanglement forging [15] is an example of the latter approach. However, a challenge faced by this framework is to find which terms and corresponding weights are more relevant in the entanglement-forging decomposition. Recently, a solution was proposed using a generative neural network [16]. Instead, in this work we argue in favour of a physically-motivated approach. In particular, we aim to exploit information about the entanglement structure of the target system [17] or, more generally, the entropy of its subsystems. These properties can indicate how to apply the decomposition and even guide the weight distribution of the product states.
We call this novel approach Entropy-Driven Entanglement Forging (EDEF). Figure 1 illustrates the EDEF algorithm, which can be applied to a physical system (middle panel) that, once encoded into qubits, can be solved with a VQA (left panel). When two low-entanglement subsystems $A$ and $B$ can be identified, they define a bipartition of the system where entanglement forging can be applied efficiently (right panel). In this alternative approach, the VQA is simplified into smaller circuits with fewer qubits that output local states on each subsystem after optimization. Finally, we recover the ground state of the entire system with a linear combination of the local states.
@@
We showcase our proposal studying two many-body systems. First, we describe the one-dimensional (1D) Fermi-Hubbard (FH) model. Here we explore the performance of the EDEF algorithm across a wide range of values for the different parameters of the Hamiltonian, which lets us identify partitions with different levels of entropy. Second, we simulate the ground states of atomic nuclei using the nuclear shell model (NSM). This approach has already been the subject of several works using quantum computing tools [18, 19, 20, 21, 22, 23, 24, 16], which point out the need for a significant amount of quantum-computing resources to simulate medium-mass nuclei [25, 24]. For both systems, we implement EDEF onto an Adaptive Derivative-Assembled Pseudo-Trotter ansatz-Variational Quantum Eigensolver (ADAPT-VQE), which has previously shown to be effective both for the FH model [26] and the NSM [18, 7]. Nonetheless, our approach can be integrated with other VQAs in a straightforward way. As illustrated in Fig. 1, we expect that the EDEF algorithm allows us to reduce both the number of qubits and the depth of the circuits of our FH and NSM many-body simulations.
@@ II.1 Entropy-driven entanglement forging
Quantum systems that are fully separable on a bipartition only need one state for each part, therefore, entanglement forging is simplest for these systems. For low entanglement between the two parts, only a few instances $\chi_{cut}$ of each subsystem are necessary to simulate it accurately, since there is a tail of terms with very small coefficients $\lambda_{i}$ that contribute negligibly to the fidelity. In contrast, for strongly-entangled subsystems one needs exponentially many states with the smallest number of qubits in one of the subsystems. Thus, in a general setting of entanglement forging, one must find a favourable bipartition, decide how many subcircuits to run, and optimize the coefficients for each of them to recover the full quantum state. Since the number of possible bipartitions in a system scales exponentially with the number of qubits $N_{q}$, entanglement forging is best suited when it is physically driven, for instance when knowledge of the system indicates low entanglement across specific sectors.
In this work, we illustrate EDEF by studying the 1D FH model with a tunable hopping term between the two central sites. This model allows us to test different levels of entropy and their impact on the quality of the results after applying one layer of EDEF. In addition, the low entanglement present between the protons and neutrons of atomic nuclei studied with the NSM [30, 31] provides an ideal practical testbed for EDEF. While we focus on the ADAPT-VQE algorithm, other VQAs can also be applied to the local circuits.
@@
6:         Set cutoff number of product states $\chi_{cut}$
@@
In this pseudocode, $A$, $B$ denote the two partitions of the system, the unitaries $\mathcal{U}^{i}_{A}$, $\mathcal{U}^{i}_{B}$ are applied individually to circuit $i$ of the decomposition on the corresponding partition, and $\varepsilon^{\prime}$ is the ground-state energy found with EDEF. The bases in the fourth step must have well-defined quantum numbers according to the model Hamiltonian and the partition. Throughout this work, we refer to the application of $l$ layers of EDEF as $l$-cut EDEF.
@@ II.2 Fermi-Hubbard model
The Fermi-Hubbard model describes fermions on a lattice and serves as a simplified model to simulate valence electrons on a crystal and fermionic ultracold gases in optical lattices [32]. In its simplest 1D form, the FH Hamiltonian includes a hopping term, $t$, accounting for the tunneling of fermions between adjacent sites, and an interaction term, $U$, which adds energy whenever a spin-up and a spin-down fermion occupy the same site. Here we consider a 1D lattice with an even number of sites, labeled $i=1,\cdots,N_{s}$, and a tunable hopping, $t_{m}$, between the two middle sites: $i=N_{s}/2,N_{s}/2+1$. The Hamiltonian reads,
@@
where $a_{j_{m}\sigma}$ ($a^{\dagger}_{j_{m}\sigma}$) are the annihilation (creation) operators for a fermion at site $j$ and spin $\sigma$, with $j=1,\cdots,N_{s}$ and $\sigma=\uparrow,\downarrow$; $n_{i\sigma}=a^{\dagger}_{j_{m}\sigma}a_{j_{m}\sigma}$ is the number operator, while $\langle i,j\rangle$ indicates pairs of first-neighbor sites, and $i_{m}=N_{s}/2$, $j_{m}=i_{m}+1$ the two middle sites. We consider repulsive interactions, $U>0$, with $t>0$, $t_{m}>0$. Here, we set $t$ as our energy unit and vary $t_{m}$ and $U$, while fixing the number of particles for each spin, $N_{\sigma}$.
Figure 2: Normalized entropy $S/S_{max}$ (black solid line) and infidelities $I_{n}$ corresponding to normalized Schmidt decompositions with $n=1,\cdots,8$ product states (red dashed lines, from top to bottom) as a function of the hopping $t_{m}$ between the two middle sites for a FH model with four sites and $t=U$.
The regular FH Hamiltonian corresponds to $t_{m}=t$, while smaller (larger) values of $t_{m}$ couple more weakly (strongly) the left and right parts of the lattice, thus reducing (enhancing) their entanglement. In a standard Jordan-Wigner mapping [33] with $N_{q}=2N_{s}$ qubits where odd ones ($i=1,\cdots,2N_{s}-1$) correspond to spin-up sites and even qubits ($i=2,\cdots,2N_{s}$) to spin-down sites, the entanglement is given by the von Neumann entropy in Eq. (1), $S$, between the first and second halves of qubits. Following Eq. (3), the entropy is upper bounded by $S_{max}=N_{s}$. Figure 2 shows $S/S_{max}$ as a function of $t_{m}$ for a lattice with $N_{s}=4$ sites, $U=t$, and half-filling, $N_{\uparrow}=N_{\downarrow}=2$. The entropy at $t_{m}=0$ vanishes as the Hamiltonian can be written as a tensor product of two parts involving only the left or right $N_{s}/2$ sites. As $t_{m}$ increases, the entropy grows up to $S/S_{max}=0.60$ at $t_{m}=2t$. For larger $t_{m}$ (not shown in Fig. 2), the entropy keeps growing and then slightly decreases, converging to $S/S_{max}=0.75$ as $t_{m}\to\infty$.
@@
$I_{n}$ always decreases for larger $n$, as the singular values $\lambda_{i}^{2}$ are non-negative real numbers sorted in decreasing order. Figure 2 also shows the infidelities for $n=1,\cdots,8$ (dashed red lines) as a function of $t_{m}$, for the FH model described in the above paragraph. For $t_{m}=0$, there is only one singular value $\lambda_{1}=1$ and all $I_{n}=0$. As $t_{m}$ increases, the $I_{n}$’s grow at different rates. This indicates that depending on the target infidelity and $t_{m}$ value, a different cutoff $n$ will be needed, and if $I_{n}$ is good enough for our purpose, then EDEF with $\chi_{cut}=n$ provides a suitable approximation. Nonetheless, the large gap between $I_{4}$ and $I_{5}$ for middle hoppings $t_{m}\lesssim t$, makes $n=5$ a particularly efficient cutoff.
@@
Figure 3: First eight singular values, $\lambda_{i}$, of the Schmidt decomposition into left and right sites of the ground state of the FH model with 4 particles, for different interaction strengths, $U$, and central hopping terms, $t_{m}$, labeled as ($t_{m}$, $U$). Empty (solid) symbols indicate $U=t$ ($U=3t$) singular values. As the central hopping is reduced, the values $i\geq 2$ get smaller.
Figure 4: Partitions used in the first and second layers of EDEF (labelled as 1st and 2nd cut) for the nuclear shell model in the $sd$ and $pf$ valence spaces. The number on top of each single-particle state labels the qubit for the implementation in ADAPT-VQE under a Jordan–Wigner mapping. The first cut always separates the proton and neutron sectors, while the second cut aims to separate orbitals according to their energy (orbitals are ordered by their energy, with lowest-energy ones at the bottom).
In the following, we fix the number of sites to $N_{s}=4$ and use the same set of central hopping terms $t_{m}$ and interactions $U$ of Fig. 3 to illustrate the performance of EDEF in different settings. Nonetheless, the distribution of spin-up and spin-down fermions discussed in this section, $N_{\uparrow}$ and $N_{\downarrow}$, generalizes to lattices with more sites (eg $N_{s}=6$, $N_{s}=8$) and particles, as long as interactions $U$ are not large (eg $U=t$) and $N_{\uparrow}=N_{\downarrow}$ is even. In these cases, the first product state consists of $N_{\uparrow}/2$ and $N_{\downarrow}/2$ in each side, while the following four product states are four-fold degenerate, corresponding to $N_{\uparrow}/2-1$ and $N_{\downarrow}/2+1$ on the left, and the corresponding degenerate product states obtained by left/right and spin-up/spin-down exchanges.
@@ II.3 Nuclear shell model
The nuclear shell model (NSM), or configuration-interaction method, is one of the most successful frameworks to study nuclear structure [34, 35, 36, 37, 38, 39]. Much alike its atomic counterpart, the NSM characterizes nuclear dynamics in a restricted configuration space, also called valence space, where nucleons effectively interact. The valence space is bounded by single-particle states which, if completely filled with nucleons, lead to magic numbers that characterize especially stable configurations associated with large energy gaps. As the nuclear interaction is rotationally invariant and nucleons are fermions, the single-particle basis states are labelled by the quantum numbers $pl_{j}$ and $m$, where $p$ is the principal quantum number, $l$ the orbital angular momentum – usually given in spectroscopic notation – and $j$ the total angular momentum with third-component projection $m$. As illustrated in Fig. 4, single-particle states with the same $pl_{j}$ and different $m$ are degenerate in energy. The effective Hamiltonian in the valence space is
@@
where $a_{i}$ ($a_{i}^{{\dagger}}$) annihilate (create) a nucleon in the single-particle state $i$ with energy $\varepsilon_{i}$. The antisymmetrized two-body matrix elements $\bar{v}_{ijkl}=v_{ijkl}-v_{ijlk}$ are obtained from the full-space nucleon-nucleon interaction. In this work, we use the standard USDB interaction [40] in the sd shell for neon, and KB3G [41] in the pf shell for titanium. Figure 4 shows the orbitals comprising these valence spaces: $0d_{5/2}$, $1s_{1/2}$ and $0d_{3/2}$ for the $sd$ shell and $0f_{7/2}$, $1p_{3/2}$, $1p_{1/2}$ and $0f_{5/2}$ for the $pf$ shell, ordered according to their single-particle energy.
@@
$|JMTT_{z}\rangle=\sum_{\alpha}a_{\alpha}|\alpha;MT_{z}\rangle,$ (8)
@@
where $M$ is the projection of the total angular momentum of the nucleus, $J$, and $T_{z}=(N-Z)/2$ the one of the total nuclear isospin, $T$. The coefficients $a_{\alpha}$ are obtained by diagonalizing the Hamiltonian in the many-body basis and guarantee that nuclear states have good $J$ and $T$ quantum numbers. State-of-the-art nuclear shell-model codes  [42, 43, 44, 45] face the challenge to build and diagonalize the Hamiltonian matrix for heavy nuclei. This is because the many-body basis, formed by the exponentially-growing set of all possible configurations for protons and neutrons in the valence space, becomes untractable for classical computers when nucleons fill about half the valence-space single-particle states.
@@
In the advent of quantum computers, alternative quantum algorithms have been proposed to find ground states within the NSM [20, 19, 23, 25, 22, 21, 16, 18, 24]. In parallel, recent works provide nuclear structure insights by analyzing nuclei in terms of quantum information measures such as the von Neumann entropy [46, 47, 48, 49, 20, 50]. Notably, among all possible bipartitions, separating proton and neutron orbitals has the lowest von Neumann entropy [31], and proton-neutron entanglement decreases further for more neutron-rich systems [30]. In fact, these properties have been used to improve nuclear-structure calculations using classical methods [51, 52].
@@
Here we apply these insights into the Schmidt decomposition of Eq. (2). Figure 5 shows the first eight singular values for various beryllium (calculated with the Cohen-Kurath interaction [53] in the $0p_{3/2}$, $0p_{1/2}$ valence space), neon, and titanium isotopes. In all nuclei, the singular values become smaller exponentially, especially for the most neutron-rich nuclei 28Ne and 60Ti, and consist of a first large non-degenerate value followed by five degenerate ones. This degeneracy can be associated with five product states with opposite $M$ in the proton and neutron states, denoted $|M^{(p)}\rangle\otimes|M^{(n)}\rangle$: |±$|\pm 2\rangle\otimes|\mp 2\rangle$, |±$|\pm 1\rangle\otimes|\mp 1\rangle$, and $|0\rangle\otimes|0\rangle$. These results suggests that nuclear ground states, especially for neutron-rich nuclei, can be well approximated within six separate product states of only protons and neutrons. Due to degeneracy and normalization, just a single coefficient is needed.
@@
Figure 5: First eight singular values, $\lambda_{i}$, of the Schmidt decomposition of into protons and neutrons for the NSM ground state of various beryllium (p shell), neon (sd shell) and titanium (pf shell) nuclei. In all cases a first large singular value is followed by much smaller five-fold degenerate second to seventh singular values, specially for neutron-rich isotopes.

### v2

@@ Abstract
Simulating physical systems with variational quantum algorithms is a well-studied approach, but it is challenging to implement in current devices due to demands in qubit number and circuit depth. We show how limited knowledge of the system, namely the entropy of its subsystems, its entanglement structure or certain symmetries, can be used to reduce the cost of these algorithms with entanglement forging. To do so, we simulate a Fermi-Hubbard one-dimensional chain with a parametrized hopping term, as well as atomic nuclei 28Ne and 60Ti with the nuclear shell model. Using an adaptive variational quantum eigensolver we find significant reductions in both the maximum number of qubits (up to one fourth) and the amount of two-qubit gates (over an order of magnitude) required in the quantum circuits. Our findings indicate that our method, entropy-driven entanglement forging, can be used to adjust quantum simulations to the limitations of noisy intermediate-scale quantum devices.
@@ I Introduction
In modern quantum devices, the number of available qubits and low-error quantum gates imposes a strong limitation in the accuracy of the final results not only in fault-tolerant schemes but also in current noisy simulations. Since quantum algorithms are applied to complex quantum many-body problems [1, 2], including quantum chemistry [3, 4, 5], condensed matter  [6, 7, 8] and nuclear physics [9, 10, 11], this constraint has been the source of many techniques that aim to simulate large systems with smaller but equivalent ones needing fewer quantum resources. Some encoding methods aim to reduce the dimension of the input data, like quantum autoencoding [12] or other physically-inspired frameworks [13]. On the other hand, other approaches focus directly on circuits to run equivalent simulations using fewer qubits. The best-known of them, circuit knitting [14, 15], takes advantage of circuits with sparsely-connected subsections in order to break them apart into smaller ones. More recently, an alternative approach has also been explored: one can train small circuits to prepare local states, and then recover the global solution with post-processing. Both types of techniques are commonly used on variational quantum algorithms (VQA) [1, 2].
Entanglement forging [16] is an example of the latter approach. However, a challenge faced by this framework is to find which terms and corresponding weights are more relevant in the entanglement-forging decomposition. Recently, a solution was proposed using a generative neural network [17]. Instead, in this work we argue in favour of a physically-motivated approach. In particular, we aim to exploit information about the entanglement structure of the target system [18] or, more generally, the entropy of its subsystems and certain symmetries of the Hamiltonian. These properties can indicate how to apply the decomposition and guide the weight distribution of the product states.
We call this novel approach Entropy-Driven Entanglement Forging (EDEF). Figure 1 illustrates the EDEF algorithm, which can be applied to a physical system (middle panel) that, once encoded into qubits, can be solved with a VQA (left panel). When two low-entanglement subsystems $A$ and $B$ can be identified, they define a bipartition of the system where entanglement forging can be applied efficiently (right panel). In this alternative approach, the VQA is simplified into smaller circuits with fewer qubits that output local states for the $A$ and $B$ subsystems after optimization. Finally, we recover the ground state of the entire system with a linear combination of these local states.
@@
We showcase our proposal studying two many-body systems. First, we use the one-dimensional (1D) Fermi-Hubbard (FH) model. Variational [6, 19, 20, 21, 8] and adiabatic [7] quantum algorithms have been devised and tested to find ground states of the FH Hamiltonian. These algorithms have focused mainly on optimizing circuit depths and on simulations of small FH lattices given the current limitations on the number of qubits per chip. Here we explore the performance of the EDEF algorithm across a range of values for the different parameters of the Hamiltonian, which lets us identify partitions with varying levels of entropy. Second, we simulate the ground states of atomic nuclei using the nuclear shell model (NSM). This approach has already been the subject of several works using quantum computing tools [22, 23, 24, 25, 26, 27, 28, 17], which point out the need for a significant amount of quantum-computing resources to simulate medium-mass nuclei [29, 28]. For both systems, we implement EDEF onto an Adaptive Derivative-Assembled Pseudo-Trotter ansatz-Variational Quantum Eigensolver (ADAPT-VQE), which has previously shown to be effective both for the FH model [21] and the NSM [22, 29]. Nonetheless, our approach can be integrated with other VQAs in a straightforward way. As demonstrated in the following sections, the use of EDEF effectively reduces both the number of qubits and the circuit depth in our FH and NSM many-body simulations. This is particularly beneficial for noisy intermediate-scale quantum devices and holds promise for advancing the broader challenge of simulating quantum many-body systems.
@@ II.1 Entropy-driven entanglement forging
Quantum systems that are fully separable on a bipartition only need one state for each part, therefore, entanglement forging is simplest for these systems. For low entanglement between the two parts, only a few instances $\chi_{\rm cut}$ of each subsystem are necessary to simulate it accurately, since there is a tail of terms with very small coefficients $\lambda_{i}$ that contribute negligibly. In contrast, for strongly-entangled subsystems one needs exponentially many states with the smallest number of qubits in one of the subsystems. Thus, in a general setting of entanglement forging, one must find a favourable bipartition, decide how many subcircuits to run, and optimize the coefficients for each of them to recover the full quantum state. Since the number of possible bipartitions in a system scales exponentially with the number of qubits $N_{q}$, entanglement forging is best suited when it is physically driven, for instance when knowledge of the system indicates low entanglement across specific sectors.
In this work, we illustrate EDEF by studying the 1D FH model with a tunable hopping term between the two central sites. This model allows us to test different levels of entropy and their impact on the quality of the results after applying one layer of EDEF. In addition, the low entanglement between protons and neutrons in atomic nuclei, as demonstrated with the NSM in Refs. [33, 34], provides an ideal practical testbed for EDEF. While we focus on the ADAPT-VQE algorithm, other VQAs can also be applied to the local circuits. We therefore expect EDEF to be algorithm-agnostic.
@@
6:         Set cutoff number of product states $\chi_{\rm cut}$
@@
In this pseudocode, $A$, $B$ denote the two partitions of the system, the unitaries $\mathcal{U}^{i}_{A}$, $\mathcal{U}^{i}_{B}$ are applied individually to circuit $i$ of the decomposition on the corresponding partition, and $\varepsilon^{\prime}$ is the ground-state energy found with EDEF. The threshold to determine the stopping point depends on the context of application, and it can be motivated by a physical purpose or by the limitations of the algorithm. The bases in the fourth step must have well-defined quantum numbers according to the Hamiltonian and the partition. Throughout this work, we refer to the application of $l$ layers of EDEF as $l$-cut EDEF.
@@ II.2 Fermi-Hubbard model
The FH model describes fermions on a lattice and serves as a simplified model to simulate valence electrons on a crystal [35] and fermionic ultracold gases in optical lattices [36]. In its simplest 1D form, the FH Hamiltonian includes a hopping term, $t$, accounting for the tunneling of fermions between adjacent sites, and an interaction term, $U$, which adds energy whenever a spin-up and a spin-down fermion occupy the same site. Here we consider a 1D lattice with an even number of sites, labeled $i=1,\cdots,N_{s}$, and a tunable hopping, $t_{m}$, between the two central sites, $i_{m}=N_{s}/2$ and $j_{m}=i_{m}+1$. The Hamiltonian reads,
@@
where $a_{i\sigma}$ ($a^{\dagger}_{i\sigma}$) are the annihilation (creation) operators for a fermion at site $i$ and spin $\sigma$, with $i=1,\cdots,N_{s}$ and $\sigma=\uparrow,\downarrow$; $n_{i\sigma}=a^{\dagger}_{i\sigma}a_{i\sigma}$ is the number operator, while $\langle i,j\rangle$ indicates pairs of first-neighbor sites. We consider repulsive interactions, $U>0$, with $t>0$, $t_{m}>0$. Here, we set $t$ as our energy unit and vary $t_{m}$ and $U$, while fixing the number of particles for each spin, $N_{\sigma}$.
Figure 2: Normalized entropy $S/S_{max}$ (black solid line) and infidelities $I_{n}$ corresponding to normalized Schmidt decompositions with $n=1,\cdots,8$ product states (dashed lines, from top to bottom) as a function of the hopping $t_{m}$ between the two middle sites for a FH model with $N_{s}=4$ sites and $U=t$.
The regular FH Hamiltonian corresponds to $t_{m}=t$, while smaller (larger) values of $t_{m}$ couple more weakly (strongly) the left and right partitions of the lattice, thus reducing (enhancing) their entanglement. In a standard Jordan-Wigner mapping [37] with $N_{q}=2N_{s}$ qubits, odd qubits ($i=1,3,\cdots,2N_{s}-1$) correspond to spin-up sites and even qubits ($i=2,4,\cdots,2N_{s}$) to spin-down sites. The entanglement is then given by the von Neumann entropy in Eq. (1), $S$, between the first and second halves of qubits, and following Eq. (3) it is upper bounded by $S_{max}=N_{s}$.
Figure 2 shows $S/S_{max}$ as a function of $t_{m}$ for a lattice with $N_{s}=4$ sites, $U=t$, and half-filling, $N_{\uparrow}=N_{\downarrow}=2$. The entropy at $t_{m}=0$ vanishes as the Hamiltonian can be written as a tensor product of two parts involving only the left or right partitions with $N_{s}/2$ sites each. As $t_{m}$ increases, the entropy grows up to $S/S_{max}=0.60$ at $t_{m}=2t$. For larger $t_{m}$ (not shown in Fig. 2), the entropy keeps growing and then slightly decreases, converging to $S/S_{max}=0.75$ as $t_{m}\to\infty$.
@@
This infidelity allows us to quantify the similarity between a state that has been cut at the $n$-th singular value and the complete, uncut state. It thus provides a useful proxy for the quality of the Schmidt decomposition of a given state into partitions. If only a few states from each partition are relevant, the infidelity should tend to zero rapidly as a function of $n$. In fact, $I_{n}$ always decreases as $n$ increases, since the singular values $\lambda_{i}^{2}$ are non-negative real numbers sorted in decreasing order. The rate of decrease as a function of $n$ is expected to depend on the entanglement structure of the system.
Figure 2 shows the infidelities for $n=1,\cdots,8$ (dashed lines) as a function of $t_{m}$, for the FH model described above. For $t_{m}=0$, there is only one singular value, so that $\lambda_{1}=1$ and $I_{n}=0$ for all values of $n$. As $t_{m}$ increases, the different infidelities $I_{n}$ grow at different rates. This indicates that, depending on the target infidelity and $t_{m}$ value, a different cutoff $\chi_{\rm cut}=n$ is needed to achieve a good quality description of the complete state $|\psi_{\chi}\rangle$. If a given value of $I_{n}$ is good enough for our purpose, the figure indicates which EDEF with $n=\chi_{\rm cut}$ provides a suitable approximation. For instance, in the physical case where $t_{m}=1$, a $1\%$ infidelity is reached with $\chi_{\rm cut}=5$. We note, in fact, that there is a large gap between $I_{4}$ and $I_{5}$ for values of the central hopping $t_{m}\lesssim t$. In fact, this gap guarantees that for $\chi_{\rm cut}=5$ the infidelity is lower than $10\%$ across all values of $t_{m}$.
@@
Figure 3: First eight singular values, $\lambda_{i}$, of the Schmidt decomposition into left and right partitions of the ground state of the FH model with 4 particles, for different interaction strengths, $U$, and central hopping terms, $t_{m}$, labeled ($t_{m}$, $U$). Empty (solid) symbols indicate singular values for an interaction strength $U=t$ ($U=3t$). As the central hopping is reduced, the values $i\geq 2$ get smaller.
Figure 2 indicates that the appearance of a gap in the infidelity is more prominent as $t_{m}$ decreases. This is somewhat natural, in that a weaker central hopping leads to a lower entanglement between left and right partitions and makes the $n=5$ cutoff more efficient. This is also clearly illustrated in the Schmidt coefficients shown in Fig. 3: the lower the value of $t_{m}$, the smaller the Schmidt coefficients are with $i\geq 2$. While the infidelity results in Fig. 2 are shown for $U=t$, different values of $U$ do not change notably the picture, as can be already anticipated from the structure of the singular values shown in Fig. 3. We find indeed that there is only a small difference in the Schmidt coefficients of systems with different values of $U$, corresponding to either empty ($U=t$) or solid ($U=3t$) symbols.
Figure 4: Partitions used in the first and second layers of EDEF (labelled as 1st and 2nd cut) for the nuclear shell model in the $sd$ (bottom panels) and $pf$ (top panels) valence spaces. Single-particle states are ordered according to their single-particle energy, with more deeply bound states at the bottom. The number on top of each state labels the qubit in our implementation of ADAPT-VQE under a Jordan–Wigner mapping. The first EDEF cut always separates the proton and neutron sectors, while the second cut aims to separate orbitals according to their single-particle energies.
In the following, we fix the number of sites to $N_{s}=4$ and use the same set of central hopping terms $t_{m}$ and interactions $U$ of Fig. 3 to illustrate the performance of EDEF in different settings. Nonetheless, the distribution of spin-up and spin-down fermions discussed in this section, $N_{\uparrow}$ and $N_{\downarrow}$, can be generalized to lattices with more sites (eg $N_{s}=6$, $N_{s}=8$, etc) and particles, as long as interactions $U$ are not large (eg $U=t$) and $N_{\uparrow}=N_{\downarrow}$ is even. In these cases, the Schmidt decomposition has the same structure as the one presented here. A first product state consists of $N_{\uparrow}/2$ and $N_{\downarrow}/2$ spins in each side. The following four product states are four-fold degenerate, corresponding to $N_{\uparrow}/2-1$ and $N_{\downarrow}/2+1$ on the left, and the corresponding degenerate product states obtained by left/right and spin-up/spin-down exchanges. In this sense, we expect the conclusions that we draw with our setup to be relatively general.
@@ II.3 Nuclear shell model
The nuclear shell model (NSM), or configuration-interaction method, is one of the most successful frameworks to study nuclear structure [38, 39, 40, 41, 42, 43]. Much alike its atomic counterpart, the NSM characterizes nuclear dynamics in a restricted configuration space, also called valence space, where nucleons effectively interact. The valence space is bounded by single-particle states which, if completely filled with nucleons, lead to magic numbers that characterize especially stable configurations associated with large single-particle energy gaps. As the nuclear interaction is rotationally invariant and nucleons are fermions, the single-particle basis states are labelled by the quantum numbers $pl_{j}$ and $m$, where $p$ is the principal quantum number; $l$, the orbital angular momentum – usually given in spectroscopic notation –; and $j$, the total angular momentum with third-component projection $m$. As illustrated in Fig. 4, single-particle states with the same $pl_{j}$ and different $m$ are degenerate in energy.
The effective Hamiltonian in the valence space is
@@
where the operators $a_{i}$ ($a_{i}^{{\dagger}}$) annihilate (create) a nucleon in the single-particle state $i$ with energy $\varepsilon_{i}$. The antisymmetrized two-body matrix elements $\bar{v}_{ijkl}=v_{ijkl}-v_{ijlk}$ can be obtained from the full-space nucleon-nucleon interaction, but are customarily fit to specific nuclear shells. In this work, we use the standard USDB interaction [44] in the sd shell for neon, and KB3G [45] in the pf shell for titanium. Figure 4 shows the orbitals comprising these valence spaces. The bottom panels show the $0d_{5/2}$, $1s_{1/2}$ and $0d_{3/2}$ states for the $sd$ shell for protons (left) and neutrons (right). The top panels display the $0f_{7/2}$, $1p_{3/2}$, $1p_{1/2}$ and $0f_{5/2}$ states for the $pf$ shell. The vertical spacing provides an indication of the different $\varepsilon_{i}$ values. The numbers correspond to the Jordan-Wigner mapping of the different states [29, 34].
@@
$|JJ_{z}TT_{z}\rangle=\sum_{\alpha}c_{\alpha}|\alpha;J_{z}T_{z}\rangle,$ (8)
@@
where $J_{z}$ is the projection of the total angular momentum of the nucleus, $J$, and $T_{z}=(N-Z)/2$ is the projection of the total nuclear isospin, $T$. The coefficients $c_{\alpha}$ are obtained by diagonalizing the Hamiltonian in the many-body basis and guarantee that nuclear states have good $J$ and $T$ quantum numbers. State-of-the-art nuclear shell-model codes  [46, 47, 48, 49] face a significant challenge to build and diagonalize the Hamiltonian matrix for heavy nuclei. This is because the many-body basis, formed by the exponentially-growing set of all possible configurations for protons and neutrons in the valence space, becomes untractable for classical computers when nucleons fill about half the valence-space single-particle states. While this is not an issue in small configurations spaces like the $p$ or the $sd$ shells, current classical simulations are significantly limited in regions of the nuclide chart with higher mass numbers.
@@
With the advent of quantum computers, alternative quantum algorithms have been proposed to find ground states within a NSM framework [24, 23, 27, 29, 26, 25, 17, 22, 28]. In parallel, recent works have provided relevant nuclear structure insights by analyzing nuclei in terms of quantum information measures such as the von Neumann entropy [50, 51, 52, 53, 24, 54] or the quantum magic [55] . A remarkable finding associated to this line of research is the recent finding that, among all possible bipartitions, the separation of a nuclear model space into proton and neutron orbitals has the lowest von Neumann entropy [34]. Morevoer, proton-neutron entanglement decreases for more neutron-rich systems [33]. These properties have been used to improve nuclear-structure calculations using classical methods [56, 57].
@@
Here, we apply these insights into the Schmidt decomposition of Eq. (2) with the aim of improving quantum simulations of nuclear structure. Figure 5 shows the first eight singular values for various beryllium  [^1] We calculate beryllium isotopes employing the Cohen-Kurath interaction [71] in the $p-$shell. We do not shown the equivalent to Fig. 4 in this configuration space for brevity., neon, and titanium isotopes. In all nuclei, the singular values become smaller exponentially, especially for the most neutron-rich isotopes 28Ne and 60Ti. The Schmidt decomposition consists of a first large non-degenerate value, with $\lambda_{1}$ close to unity, followed by five degenerate vectors with coefficients $10^{-2}<\lambda_{2-6}^{2}<10^{-1}$. This picture is qualitatively similar to that found in the FH model in Fig. 3, except that the second set of states has $5$ rather than $4$ degenerate states.
The degeneracy in the NSM can be associated with five product states with opposite $M$ in the proton and neutron states, which we denote $|J_{z}^{(p)}\rangle\otimes|J_{z}^{(n)}\rangle$: |±$|\pm 2\rangle\otimes|\mp 2\rangle$, |±$|\pm 1\rangle\otimes|\mp 1\rangle$, and $|0\rangle\otimes|0\rangle$. These results suggests that nuclear ground states, especially for neutron-rich isotopes, can be well approximated by six separate product states of only protons and neutrons. As a consequence, and due to this degeneracy and the overall normalization of the state, a single coefficient is needed to describe the full state in terms of very few Schmidt vectors.
@@
Figure 5: First eight singular values, $\lambda_{i}$, of the Schmidt decomposition into protons and neutrons for the NSM ground state of different beryllium (p shell, solid circles and squares), neon (sd shell, empty symbols) and titanium (pf shell, solid diamond and triangle) nuclei. In all cases, a first large singular value is followed by much smaller, five-fold degenerate second to seventh singular values. The Schmidt coefficients for these degenerate singular values are smaller in neutron-rich isotopes.
The Schmidt coefficients beyond $i>6$ in Fig. 5 are relatively small compared to the first $6$ eigenvalues. For the Ne and Ti isotopes shown in the figure, we find that as the systems become more neutron rich, the corresponding Schmidt coefficients for $i>6$ become smaller. Whereas $\lambda_{7}^{2}\approx 10^{-2}$ for 20Ne, the same coefficient for 28Ne is two orders of magnitude smaller, $\lambda_{7}^{2}\approx 10^{-4}$. A similar pattern is observed in titanium, between 44Ti and 60Ti. In the following, we choose 28Ne and 60Ti as our reference isotopes for the EDEF analysis, since we expect these to be best suited for the approach.